# {adjustwidth}0cm-0.2cm Higgs-Induced Gravitational Waves: the Interplay of Non-Minimal Couplings, Kination and Top Quark Mass

### v1

@@ 2 Effective Higgs potential with non-minimal curvature coupling
The transition from an inflationary epoch to a kination-dominated phase introduces a significant tachyonic mass in the Higgs effective potential, triggering a second-order phase transition at high energy scales. This tachyonic instability can drive the Higgs fluctuations towards the unstable region of the effective potential as they become exponentially amplified. At high energies, the zero-temperature effective potential is approximately given by
@@
where $h$ is the Higgs degree of freedom in the unitary gauge $H=(0,\,h/\sqrt{2})^{T}$. Notice that the energy scale $\mu$ entering the Renormalisation-Group-Improved (RGI) running $\lambda(\mu)$ is set as a combination of the background cosmological scale and the amplitude of the Higgs field itself $\mu=\mathcal{H}+h$, as per the standard prescription on curved backgrounds Markkanen:2018bfx ; Markkanen:2017dlc . To facilitate the analytical and numerical treatments, we make use of the following parametrisation Bezrukov:2014bra of the three-loop $\overline{\rm MS}$ running of the Higgs self-coupling in the vicinity of the instability scale,
$\lambda^{\rm 3loop}(\mu)=\lambda_{0}+b\log^{2}\left[\frac{\mu}{q\,\rm M_{P}}\right]\;,$ (3)
which has an explicit dependence on the Higgs and top quark masses via the parameters $\lambda_{0}$, $b$ and $q$, namely
$\lambda_{0}=0.003297\left(\left(\frac{m_{h}}{\textrm{GeV}}-126.13\right)-2\left(\frac{m_{t}}{\textrm{GeV}}-171.5\right)\right)\;,$
$q=0.3\exp\left[0.5\left(\frac{m_{h}}{\textrm{GeV}}-126.13\right)-0.03\left(\frac{m_{t}}{\textrm{GeV}}-171.5\right)\right]\;,$
$b=0.00002292-1.12524\times 10^{-6}\left(\left(\frac{m_{h}}{\textrm{GeV}}-126.13\right)-1.75912\left(\frac{m_{t}}{\textrm{GeV}}-171.5\right)\right)\;.$ (4)
@@
We will assume from now on a fixed value of the Higgs mass ${m_{h}=125.2\textrm{ GeV}}$ ParticleDataGroup:2024cfk and focus on the impact the top quark mass has on the running of $\lambda$. A graphical representation of the effective potential can be seen in Figure 1, where the green and red lines exemplify the possible formation of a barrier in the SM effective potential at the instability scale $\mu_{\rm inst}$ for different choices of top quark pole mass.
Imposing a stability constraint means ensuring that the tachyonically amplified fluctuations do not exceed the instability scale. The parametric description computed in a previous work Laverda:2023uqv provides a convenient way to estimate the Higgs energy density at the end of the first semi-oscillation. This can be compared to the height of the barrier in the effective potential, setting with it a stability constraint, namely
@@
where $\rho_{\rm tac}$ is the Higgs energy density at the end of the tachyonic phase (first semi-oscillation of the field) and $h_{\rm min}$, $h_{\rm max}$ indicate the positions of the new Hubble-induced minimum and the barrier. The tachyonically-enhanced energy density was found to follow the parametric behaviour
$\rho_{\text{tac}}(\mathcal{H}_{\rm kin},\nu,m_{t})=16\,\mathcal{H}^{4}_{\rm kin}\,\exp\left(\beta_{1}(m_{t})+\beta_{2}(m_{t})\,\nu+\beta_{3}(m_{t})\log\nu\right)\,,$ (6)
@@
obtained from 3+1-dimensional classical numerical simulations performed with the code $\mathcal{C}osmo\mathcal{L}attice$ Figueroa:2021yhd ; Figueroa:2016wxr . The coefficients $\beta_{1},\beta_{2}$ and $\beta_{3}$ in this expression are given in terms of $\nu=\sqrt{3\xi/2}$ and $n=-\log_{10}(\lambda(\mu))$, but the running in (3) allows us to trade the dependence on $\lambda$ for a dependence on $m_{t}$, thus making explicit the connection between cosmological and SM parameters. The quantity $\mathcal{H}_{\rm kin}$ is the scale corresponding to the beginning of the kination phase which, for an instantaneous transition like the one under consideration,  [^4] One might wonder if taking into account the non-instantaneous transition from inflation to kination can lead to a non-negligible contribution to the Higgs energy density. This is important whenever the tachyonic process is inefficient, i.e. when the non-minimal-coupling induced vacuum is displaced at small field values. In general, a scalar field with a time-varying effective mass between inflation and kination undergoes a process of non-adiabatic particle production. This can lead to some initial excitation that can contribute to the final energy-density after the Hubble-induced phase transition. One can estimate the energy-density produced by such process following the reference Mukhanov:2007zz and find it to be proportional to $\mathcal{H}^{4}_{\rm kin}\sqrt{\xi/12}$. For Higgs-like values of $\lambda(\mu)$, one concludes that the tachyonic production is more effective by several orders of magnitude. Modelling the non-instantaneous onset of kination has been considered in Figueroa:2024asq by including an inflaton field together with a specific potential that defines a model-dependent transition. can be identified with the inflationary scale $\mathcal{H}_{\rm kin}\approx\mathcal{H}_{\rm inf}$. A detailed discussion of the stability/instability scenarios was carried out in Laverda:2024qjt , alongside the study of the reheating stage in the Higgs, gauge and fermionic sectors, concluding than only the former is relevant for the macroscopic dynamics of heating. [^5] Resonant production of gauge bosons occurs as the Higgs oscillates in its effective potential. While this process involves successive exponential amplification, its efficiency is limited compared to the rapid tachyonic production of the Higgs itself. Assuming the Higgs oscillates uniformly in a quartic potential, resonant amplification proceeds with a Floquet index of $\mu_{k}\to\mu_{\text{max}}=0.2377$ for $q=g^{2}/\lambda\gg 1$, typical for SM couplings Greene:1997fu . Even under optimal conditions, it takes at least 10 oscillations for the energy density of bosonic fields to match that of the Higgs. The additional decay of gauge bosons into fermions does not significantly impact the heating timeline or vacuum stability constraints. However, non-Abelian interactions may aid in the thermalisation of the SM plasma, accelerating the approach to equipartition through boson scatterings and annihilations Enqvist:2015sua ; Bodeker:2007fw ; Bezrukov:2014ipa ; Repond:2016sol . Here we only emphasise that the same set of parametric formulas derived in Laverda:2023uqv allows us to quantify the course of the heating stage. Indeed, the first consequence of the tachyonic instability is the enhancement of the Higgs energy density and the reheating of the primordial Universe. To describe this phase, we make use of the radiation timescale at which the equation-of-state parameter for the Higgs field becomes $w_{h}=1/3$,
@@
evaluated at $z_{\rm rad}$, where $z=a_{\rm kin}\sqrt{6\xi}\mathcal{H}_{\rm kin}\tau$ is a rescaled conformal-time variable. The full expressions, including the numerical coefficients, can be found in Appendix A. The duration of the heating phase is encoded in the heating efficiency of the Higgs sector Rubio:2015zia ,
@@
[…]
@@
The successful heating of the Universe depends mostly on the scale of kination and on the non-minimal coupling parameter $\nu$, making the threshold scale $\mathcal{H}_{\rm kin}\simeq 10^{5.5}\text{ GeV}$ independent from the chosen value of $m_{t}$. Indeed, for energy scales far below the instability scale, the running of the self-coupling converges to $\lambda(\mu)\simeq 0.1$ for all top masses, and only a weak dependence on $\nu$ remains for small values of non-minimal interactions.
@@ 3 Stochastic gravitational-wave background
The second major consequence of the initial amplification of the Higgs quantum modes is the generation of large fluctuations on sub-horizon scales. These “bubbly” features in the field distribution lead to the appearance of large spatial gradients Laverda:2023uqv , thus sourcing a strong GW signal GarciaBellido:2008ab . This production phase is constrained to the interval between backreaction and the radiation timescale $z_{\rm rad}$, during which the non-minimal and self-interactions have comparable magnitudes. By the radiation time, the system has quenched into a state of fragmented, small fluctuations. Hence, the GW spectrum possesses a clear peaked shape, as is typical for transient sources. The features of such GW signal were studied in detail with 3+1 classical lattice simulations in a recent work Bettoni:2024ixe . The general shape of the spectrum is well described by a prototypical broken-power-law shape function Lewicki:2020azd ; Lewicki:2022pdb
@@
where the momentum and amplitude at the peak are given in terms of parametric functions obtained from numerical lattice simulations,
@@
where momenta $\kappa$ are normalised with respect to $\mathcal{H}_{\rm kin}$. As noted in the previous section, the dependence on the running of the Higgs self-coupling has been recast into a dependence on the top quark mass thanks to the analytic form of the running in (3). The IR slope of the spectrum follows the expected causal behaviour $a=3.00$ Caprini:2009fx ; Cai:2019cdl , while the UV slope depends on the dynamics of the turbulent cascade following the out-of-equilibrium tachyonic particle production with typical values of the parameters $b=152.34-6.57\,\nu$ and $c=105.85-4.79\,\nu$. Spectra-related quantities are evaluated at the radiation time, i.e. at the end of the violent GW production process. Moreover, we have used the following normalisation of the GW energy density
@@
which depend on the duration of kination via the heating efficiency $\Theta_{\rm ht}^{h}$. As discussed in Bettoni:2024ixe , the typical frequency of the spectrum does not depend on the scale of kination because of the cancellation caused by the common dependence on $\mathcal{H}_{\rm kin}$ of all relevant quantities. The heating efficiency $\Theta_{\rm ht}^{h}$ and the scale factor $a_{\rm rad}$ at the radiation time can be readily evaluated using the parametric formulas in the Appendix A.
4 Probing the top quark mass with gravitational waves
Figure 2: Energy density $\Omega_{\rm GW,0}$ and peak frequency $f_{\rm GW,0}$ of the GW spectrum at the present day as a function of $\mathcal{H}_{\rm kin}$ and $\nu$. The top quark mass has been fixed to $m_{t}=171\textrm{ GeV}$ compatible with the absolute stability of the electroweak vacuum. The coloured regions indicate the areas excluded by the bound on the inflationary scale from CMB measurements (green), by the number of effective degrees of freedom at BBN (grey), and by the minimum reheating temperature (blue).
In the most minimal scenario, the tachyonic phase in the Higgs dynamics is enough to exponentially enhance the Higgs energy density and lead to the reheating of the Universe. The details of the symmetry breaking depend on the running of the Higgs self-coupling in the chosen RGI scheme. By focussing on the dependence on the SM parameters $m_{h}$ and $m_{t}$ in (3) we can directly observe the influence that the top quark mass has on the spectrum of GWs. The specific characteristics of the GW signal would enable us to place indirect constraints to the top quark mass from gravitational physics.
The existence of an instability scale for larger values of the top quark mass forces us to consider two distinct possibilities. The scale of kination has to be a few orders of magnitude smaller than the instability scale to guarantee classical stability via the constraint

### v2

@@ 2 The spectator Higgs field approximation
The transition from an inflationary epoch to a kination-dominated phase induces a significant tachyonic mass in the Higgs effective potential, triggering a second-order phase transition at high-energy scales. To determine the limit of validity of the spectator field approximation, we examine the Einstein equations
$M_{P}^{2}\,G_{\mu\nu}=T^{(h)}_{\mu\nu}+T^{(\phi)}_{\mu\nu}\,$, (6)
with $T^{(\phi)}_{\mu\nu}$ the inflaton energy-momentum tensor. Taking the trace of this expression, performing a spatial average of both sides of the result, and using (5) to solve explicitly for $R$, we obtain
$M_{P}^{2}\,R=\langle T^{(\phi)}\rangle\frac{1-\Delta_{1}}{1+\Delta_{2}}\,$, (7)
with
$\Delta_{1}=(6\xi-1)\frac{\langle T_{\xi=0}^{(h)}\rangle}{\langle T^{(\phi)}\rangle}\,,\quad\quad\quad\Delta_{2}=(6\xi-1)\frac{\xi\langle h^{2}\rangle}{M_{P}^{2}}\,$, (8)
and $\langle T^{(\phi)}\rangle$ and $\langle T_{\xi=0}^{(h)}\rangle$ the average traces of the inflaton energy-momentum tensor and the Higgs energy-momentum tensor (5) at zero non-minimal coupling. In the conformal limit ($\xi=1/6$) and whenever the corrections $\Delta_{1}$ and $\Delta_{2}$ are negligible, the evolution of the Ricci scalar is entirely determined by the homogeneous inflaton component. These two corrections, however, play distinct roles. The term $\Delta_{1}$ controls the ratio between the $\xi$-enhanced energy density of the spectator field and that of the inflaton, thereby affecting potential sign changes in the Ricci scalar and ultimately governing the termination of the tachyonic instability. For sufficiently small self-couplings, this condition is equivalent to imposing the constraint $\langle h^{2}\rangle^{1/2}\ll M_{P}/\sqrt{\xi}$ Figueroa:2024asq , in full agreement with the assumptions in previous studies Bettoni:2019dcw ; Laverda:2023uqv ; Bettoni:2021zhq . Meanwhile, $\Delta_{2}$ quantifies the degree of non-linearity introduced by the non-minimal coupling of the Higgs field to gravity, in contrast to its behaviour in the minimally coupled Standard Model and General Relativity. As shown in Appendix A, the condition $\Delta_{2}\ll 1$ ensures that the typical fluctuations of the Higgs field remain well below the quantum cut-off scale $\Lambda_{\xi}=M_{P}/\xi$ of this inherently non-renormalizable framework Barbon:2009ya ; Burgess:2009ea ; Hertzberg:2010dc ; Burgess:2010zq ; Bezrukov:2010jz ; Bezrukov:2012hx ; Ren:2014sya ; Bezrukov:2014ipa ; Fumagalli:2017cdo ; Antoniadis:2021axu ; Mikura:2021clt ; Ito:2021ssc ; Escriva:2016cwl ; Karananas:2022byw . Beyond this threshold, and in accordance with the standard effective field theory logic, the starting Lagrangian density (1) must necessarily be supplemented with an infinite series of higher-dimensional operators to maintain consistency, regardless of the presence or absence of backreaction on the inflaton field.
3 Electroweak vacuum stability and the onset of heating
Due to the absence of a global minimum, runaway inflationary potentials like the one under consideration do not lead to standard reheating scenarios based on non-perturbative resonant production. On top of that, the suppression of tree-level interactions between the Higgs and the inflaton fields in the Einstein frame does not allow for reheating mechanisms à la instant reheating Felder:1998vq . Indeed, within the validity regime of the graviscalar effective field theory, a possible interaction portal between the inflaton and the Higgs field is not active, since the Einstein-frame Lagrangian contains only Planck-suppressed tree-level interactions between the two, see Appendix A. We also assume that, if a direct tree-level coupling of the form $\sim g^{2}h^{2}\phi^{2}$ is present in the Jordan frame—explicitly breaking the advocated shift symmetry of the inflationary potential near the crossover scale—it is suppressed by a small coupling $g$.  [^2] This coupling causes a violation of the adiabaticity condition in the spectator field and the production of particle excitations. However, the efficiency of such process is proportional to $g^{2}$ and therefore suppressed for perturbative couplings. For instance, setting $\mathcal{H}_{\rm kin}=10^{11}\textrm{ GeV}$ and $\lambda\sim 10^{-2}$, this this contribution is subdominant with respect to the tachyonic production over the whole parameter space under consideration already for $g\sim 10^{-5}$ Bettoni:2021qfs ; Laverda:2023uqv .
As a consequence, leveraging on the time-dependence of the Higgs effective mass, the tachyonic amplification of Higgs fluctuations during HIPTs offers a compelling mechanism for heating the Universe in non-oscillatory models of inflation. This amplification is typically a brief process, as both backreaction effects and the rapid decrease of the Hubble rate cause the non-minimal interaction to quickly become subdominant, effectively restoring the original symmetry of the potential Bettoni:2019dcw ; Bettoni:2021zhq ; Laverda:2023uqv . Due to the quartic shape of the SM Higgs potential, the classicalised Higgs energy density evolves asymptotically as radiation, $\rho_{h}\sim a^{-4}$, allowing it to grow over the rapidly decaying inflaton background, which, for a stiff equation of state $w\simeq w_{\phi}\simeq 1$, evolves as $\rho_{\phi}\sim a^{-6}$. The final viability of this appealing heating scenario is nonetheless deeply connected to the stability of the Standard Model vacuum at high energies. To see this, let us consider the zero-temperature Renormalisation-Group-Improved (RGI) potential for the non-minimally coupled Higgs field,
@@ 3 Electroweak vacuum stability and the onset of heating
with the scale $\mu$ in the running coupling constant $\lambda(\mu)$ set as a combination of the Higgs field itself and the cosmological background scale, $\mu^{2}=\mathcal{H}^{2}+h^{2}$, as per the standard prescription on curved cosmological backgrounds Markkanen:2018bfx ; Markkanen:2017dlc .
To facilitate the analytical and numerical treatments, we start by performing a fitting of the full numerical three-loop $\overline{\rm MS}$ running of the Higgs self coupling BezrukovNotebook parametrising its dependence on the top quark mass. Motivated by their small experimental error, we set the Higgs mass to the central value of the current best measure $m_{h}=125.20\pm 0.11\textrm{ GeV}$ and the strong coupling constant to $\alpha_{s}=0.1180$ ParticleDataGroup:2024cfk . A good ansatz in the energy-scale range $\mu\in[10^{4}\textrm{ GeV},\,10^{23}\textrm{ GeV}]$ for the parametric shape is given by
$\lambda_{\rm fit}(\mu)=\lambda_{0}+c_{1}\log\left(\frac{\mu}{\rm GeV}\right)+c_{2}\log^{2}\left(\frac{\mu}{\rm GeV}\right)+c_{3}\log\left(\log\left(\frac{\mu}{\rm GeV}\right)\right)$ (10)
where the top-mass-dependent coefficients $\lambda_{0}$, $c_{1}$, $c_{2}$ and $c_{3}$ are given by
$\lambda_{0}(m_{t})$ $=7.31965-9.13428\times 10^{-2}\times\frac{m_{t}}{\rm GeV}+2.94355\times 10^{-4}\times\left(\frac{m_{t}}{\rm GeV}\right)^{2}\;,$
$c_{1}(m_{t})$ $=1.98329\times 10^{-1}-2.34181\times 10^{-3}\times\frac{m_{t}}{\rm GeV}+7.05298\times 10^{-6}\times\left(\frac{m_{t}}{\rm GeV}\right)^{2}\;,$
$c_{2}(m_{t})$ $=-1.4835\times 10^{-4}+7.88043\times 10^{-7}\times\frac{m_{t}}{\rm GeV}\;,$
$c_{3}(m_{t})$ $=-4.19967+5.30679\times 10^{-2}\times\frac{m_{t}}{\rm GeV}-1.70896\times 10^{-4}\times\left(\frac{m_{t}}{\rm GeV}\right)^{2}\;.$ (11)
The accuracy of such parametric function is always within 1% of the full numerical running for the renormalisation scales under considerations and for a wide range of top quark masses $m_{t}\in[169\textrm{ GeV}\;,175\textrm{ GeV}]$. Note that in these expressions, $m_{t}$ refers to the top-quark pole mass $m_{t}^{\text{pole}}$, which differs from the Monte Carlo reconstructed mass $m_{t}^{\text{rec}}$ typically obtained using event generators such as PYTHIA or HERWIG Hoang:2020iah ; Nason:2017cxd . The discrepancy between these two scales is approximately 1–2 GeV, with recent measurements consistently finding $m_{t}^{\text{pole}}$ to be lower than $m_{t}^{\text{rec}}$. Specifically, $m_{t}^{\text{pole}}=170.5\pm 0.8$ GeV CMS:2019esx , while $m_{t}^{\text{rec}}=171.77\pm 0.37$ GeV CMS:2023ebf .
@@
A graphical representation of the effective potential (9) can be seen in Figure 1, where the green and red lines exemplify the possible formation of a barrier in the SM effective potential at the instability scale $\mu_{\rm inst}$ for different choices of top-quark pole mass. Beyond that scale, the Higgs potential becomes negative, with the new and significantly deeper minimum appearing at super-Planckian field values in the absence of new physics.
Crossing the electroweak vacuum instability threshold after a HIPT could trigger the formation of anti-de Sitter regions expanding at the speed of light and ultimately engulfing our observable Universe Espinosa:2015qea . To prevent this catastrophic scenario, a stability constraint must be imposed on the energy density of tachyonically amplified Higgs fluctuations. The parametric description derived in a previous study from an extensive set of numerical 3+1 classical lattice simulations Laverda:2023uqv provides a practical method for estimating the Higgs energy density when backreaction effects become relevant. Indeed, lattice simulations allow for the evolution of the interacting Higgs field in a high-occupation-number regime while fully incorporating the backreaction arising from self-interactions in the fundamental Lagrangian. By examining the complete dynamics, one can define a backreaction timescale, i.e. the moment when Higgs fluctuations reach large amplitudes and self-scattering effects become significant. The energy density at this backreaction time can then be compared to the height of the barrier in the effective potential, thereby establishing a stability constraint,
@@
with $\rho_{\rm tac}$ the Higgs energy density including the non-minimal coupling terms at the end of the tachyonic phase. The field values $h_{\rm min}$ and $h_{\rm max}$ correspond to the positions of the new Hubble-induced minimum and the top of the barrier, found by solving $\partial V_{\rm eff}/\partial h=0$. In order to obtain a good estimate of the Higgs self-coupling at the time when backreaction effects set in, we compute $\lambda(h_{\rm min})$ at the time of backreaction $z_{\rm br}\sim 10$, which corresponds to $\mathcal{H}_{\rm br}\sim 0.5\mathcal{H}_{\rm kin}$ Opferkuch:2019zbd . Indeed, the top-mass-dependent $h_{\rm min}$ scale is the typical amplitude of Higgs fluctuations a the end of the tachyonic phase.
Our numerical analysis Laverda:2023uqv using 3+1-dimensional numerical simulations with the code $\mathcal{C}osmo\mathcal{L}attice$ Figueroa:2021yhd ; Figueroa:2016wxr found the following parametric behaviour for the energy-density of the Higgs field at the end of the tachyonic phase
$\rho_{\text{tac}}(\mathcal{H}_{\rm kin},\nu,m_{t})=16\,\mathcal{H}^{4}_{\rm kin}\,\exp\left(\beta_{1}(m_{t})+\beta_{2}(m_{t})\,\nu+\beta_{3}(m_{t})\log\nu\right)\,.$ (13)
@@
The coefficients $\beta_{1},\beta_{2}$ and $\beta_{3}$ in this expression are given in terms of $n=-\log_{10}(\lambda(\mu))$, but the running in (10) allows us to trade the dependence on $\lambda$ for a dependence on $m_{t}$, thus making explicit the connection between cosmological and SM parameters. The quantity $\mathcal{H}_{\rm kin}$ in this expression is the Hubble scale corresponding to the beginning of the kination phase which, for an instantaneous transition like the one under consideration,  [^3] One might wonder if taking into account the non-instantaneous transition from inflation to kination can lead to a non-negligible contribution to the Higgs energy density. This is important whenever the tachyonic process is inefficient, i.e. when the non-minimal-coupling induced vacuum is displaced at small field values. In general, a scalar field with a time-varying effective mass between inflation and kination undergoes a process of non-adiabatic particle production. This can lead to some initial excitation that can contribute to the final energy-density after the Hubble-induced phase transition. One can estimate the amplitude reached after this non-adiabatic excitation following the references Mukhanov:2007zz ; Herranen:2015ima , and find it to be $\langle h^{2}\rangle\sim\mathcal{H}^{2}_{\rm kin}\xi$. For Higgs-like values of $\lambda(\mu)$, one concludes that the tachyonic amplification is more effective by at least one order of magnitude. Modelling the non-instantaneous onset of kination has been considered in Figueroa:2024asq by including an inflaton field together with a specific potential that defines a model-dependent transition. can be identified with the inflationary scale $\mathcal{H}_{\rm kin}\approx\mathcal{H}_{\rm inf}$. A detailed discussion of the stability/instability scenarios was carried out in Laverda:2024qjt , alongside the study of the heating stage in the Higgs, gauge and fermionic sectors, concluding than only the former is relevant for the macroscopic dynamics of heating. [^4] Resonant production of gauge bosons occurs as the Higgs oscillates in its effective potential. While this process involves successive exponential amplification, its efficiency is limited compared to the rapid tachyonic production of the Higgs itself. Assuming the Higgs oscillates uniformly in a quartic potential, resonant amplification proceeds with a Floquet index of $\mu_{k}\to\mu_{\text{max}}=0.2377$ for $q=g^{2}/\lambda\gg 1$, typical for SM couplings Greene:1997fu . Even under optimal conditions, it takes at least 10 oscillations for the energy density of bosonic fields to match that of the Higgs. The additional decay of gauge bosons into fermions does not significantly impact the heating timeline or vacuum stability constraints. However, non-Abelian interactions may aid in the thermalisation of the SM plasma, accelerating the approach to equipartition through boson scatterings and annihilations Enqvist:2015sua ; Bodeker:2007fw ; Bezrukov:2014ipa ; Repond:2016sol . Here we only emphasise that the same set of parametric formulas derived in Laverda:2023uqv allows us to quantify the course of the heating stage. Indeed, the first consequence of the Hubble-induced tachyonic instability is the enhancement of the Higgs energy density and the heating of the primordial Universe. To describe this phase, we make use of the radiation timescale at which the equation-of-state parameter for the Higgs field becomes $w_{h}=1/3$,
@@
evaluated at $z_{\rm rad}$, with $z=a_{\rm kin}\sqrt{6\xi}\mathcal{H}_{\rm kin}\tau$ a rescaled conformal-time variable. The full expressions, including the numerical coefficients, can be found in Appendix B. The duration of the heating phase is effectively encoded in the so-called heating efficiency of the Higgs sector Rubio:2015zia ,
@@
with $g_{*}^{\rm ht}=106.75$ the SM number of relativistic degrees of freedom at energies above ${\cal O}(100)\text{ GeV}$ and $\rho^{\rm ht}_{h}=\rho^{\rm ht}_{\phi}$ the total energy density of the Higgs field at the end of the heating phase. The fitting formula for $\rho_{\rm rad}$ in (15) allow us to estimate the (re)heating temperature in terms of the model parameters $(\mathcal{H}_{\rm kin},\,\nu,\,m_{t})$. Note that our definition of radiation temperature is agnostic about the achievement of a thermalised Universe. However, the thermalisation timescale is typically similar to the duration of the heating phase Micha:2004bv ; Bettoni:2021zhq and we will assume the Universe to be thermalised at the end of the heating stage.
@@
The successful heating of the Universe depends mostly on the scale of kination and on the non-minimal coupling parameter $\nu$, making the threshold scale $\mathcal{H}_{\rm kin}\simeq 10^{5.5}\text{ GeV}$ independent from the chosen value of $m_{t}$. Indeed, for energy scales far below the instability scale, the running of the Higgs self-coupling converges to $\lambda(\mu)\simeq 0.1$ for all top-quark masses, and only a weak dependence on $\nu$ remains for small values of non-minimal interactions.
@@ 4 Stochastic gravitational-wave background
The second major consequence of the initial amplification of the Higgs quantum modes is the generation of large fluctuations on sub-horizon scales. These “bubbly” features in the field distribution lead to the appearance of large spatial gradients Laverda:2023uqv , thus sourcing a strong GW signal GarciaBellido:2008ab . This production phase is limited to the interval between backreaction and the radiation timescale $z_{\rm rad}$, during which the non-minimal and self-interaction contributions have comparable magnitudes. By the radiation time, the system has quenched into a state of fragmented, small fluctuations. Hence, the GW spectrum possesses a clear peaked shape, as is typical for transient sources. The features of such GW signal were studied in detail with 3+1 classical lattice simulations in a recent work Bettoni:2024ixe . The general shape of the spectrum is well described by a prototypical broken-power-law shape function Lewicki:2020azd ; Lewicki:2022pdb
@@
where the momentum and amplitude at the peak are given in terms of parametric functions obtained from numerical 3+1 classical lattice simulations,
@@
with momenta $\kappa$ normalised with respect to $\mathcal{H}_{\rm kin}$.  [^5] Note that the typical momenta around the peak of the GW spectrum are typically $\sim 10-100\times\mathcal{H}_{\rm kin}$ and are well below the effective-field-theory cut-off scale for $\mathcal{H}_{\rm kin}\lesssim 10^{12}\textrm{ GeV}$. As noted in the previous section, the dependence on the running of the Higgs self-coupling has been recast into a dependence on the top-quark mass, thanks to the analytic form of the running in (10). The IR slope of the spectrum follows the expected causal behaviour $a=3.00$ Caprini:2009fx ; Cai:2019cdl , while the UV slope depends on the dynamics of the turbulent cascade following the out-of-equilibrium tachyonic particle production with typical values of the parameters $b=152.34-6.57\,\nu$ and $c=105.85-4.79\,\nu$. Spectra-related quantities are evaluated at the radiation time, i.e. at the end of the violent GW production process. Moreover, we have used the following normalisation of the GW energy density
@@
which depend on the duration of kination via the heating efficiency $\Theta_{\rm ht}^{h}$. As discussed in Bettoni:2024ixe , the typical frequency of the spectrum does not depend on the scale of kination because of the cancellation caused by the common dependence on $\mathcal{H}_{\rm kin}$ of all relevant quantities. The heating efficiency $\Theta_{\rm ht}^{h}$ and the scale factor $a_{\rm rad}$ at the radiation time can be readily evaluated using the parametric formulas in the Appendix B.
5 Probing the top-quark mass with gravitational waves
Figure 2: Energy density $h^{2}\Omega_{\rm GW,0}$ and peak frequency $f_{\rm GW,0}$ of the GW spectrum at the present day as a function of $\mathcal{H}_{\rm kin}$ and $\nu$. The top-quark mass has been set to be compatible with the absolute stability of the electroweak vacuum $m_{t}=170.8\textrm{ GeV}$. The coloured regions indicate the areas excluded by the proximity to the effective-theory cut-off scale (light blue), by the bound on the inflationary scale from CMB measurements (green), by the number of effective degrees of freedom at BBN (grey), and by the minimum (re)heating temperature (blue).
In the most minimal scenario, the tachyonic phase in the Higgs dynamics is enough to exponentially enhance the Higgs energy density and lead to the heating of the Universe. The details of the symmetry breaking depend on the running of the Higgs self-coupling in the chosen RGI scheme. By focussing on the dependence on the SM parameter $m_{t}$ in (10) we can directly observe the influence that the top-quark mass has on the spectrum of GWs. The specific characteristics of the GW signal enable us to place indirect constraints to the top-quark mass from gravitational physics.
The existence of an instability scale for larger values of the top-quark mass forces us to consider two distinct possibilities, since the scale of kination has to be a few orders of magnitude smaller than the instability scale to guarantee classical stability via the constraint